OpenStudy (lin.ivory):

f(x)=(3(x^2)-1)/(4-(x^2)) find horizontal and vertical asymptotes with limits

OpenStudy (anonymous):

To find the horizontal symptote(s), find the limit of the function as x ------> infinity

OpenStudy (anonymous):

As x --------> infinity, the function approaches -3, so y = -3 is a horizontal asymptote.

OpenStudy (anonymous):

For the vertical asymptotes, set denominator equal to zero, x = 2 and x = -2.

OpenStudy (lin.ivory):

i understood for the vertical aysmptotes that it was 2 and -2 by simply factoring, however when i looked at the solution, i didnt understand this! \lim_{x \rightarrow 2^{-}} f(x) =\left(\begin{matrix}11 \\ 0^{+}\end{matrix}\right) = +infty

OpenStudy (lin.ivory):

\[\lim_{x \rightarrow 2^{-}} f(x) =\frac{ 11 }{ 0^{+} } = +\infty\]

OpenStudy (lin.ivory):

\[\lim_{x \rightarrow -2^{-}} f(x) =\frac{ 11 }{ 0^{-} } = - \infty\]

OpenStudy (lin.ivory):

my question is, x is approaching from the left side of both -2 and 2, by why is one negative infinity and one is positive infinity??

OpenStudy (agent0smith):

You can plug in numbers (like -2.001 for a number left of -2, to see why, or 1.99 for a number left of 2) or graph the function to see it's behaviour near asymptotes http://www.wolframalpha.com/input/?i=%283x%5E2-1%29%2F%284-x%5E2%29

OpenStudy (lin.ivory):

ohh ok thank you!!